When the Botswana Meat Commission (BMC) increased beef price from P10 per Kg to

P16.50 per Kg, the supply of beef cattle to the BMC rose from six hundred per week to 900

per week. Use the midpoint formula to calculate the coefficient of elasticity of supply for

beef to the BMC and interpret the elasticity coefficient.


1
Expert's answer
2021-12-12T16:34:49-0500

Elasticity of supply,Es;

"=\frac{Q2-Q1}{\frac{Q2+Q1}{2}}" "\div\frac{P_1-Q_1}{\frac{P_2+Q_1}{2}}"


"=\frac{900-600}{\frac{900+600}{2}}÷\frac{16.50-10}{\frac{16.50+10}{2}}"


"=\frac{0.4}{0.491}"


"=0.82"

Since the coefficient of elasticity of supply is less than 1,the supply of the meat is referred to as inelastic.